Please show work.

14. Suppose the prices of Stock A and B change over time as follows. t+1 t+2 Pt $100 Pt $100 There are two investment scenarios Stock A Stock B +1$200 Pt+1$50 Pt+2- $100 Pt+2 -$100 a) Invest S100k in stock A and $100k in stock B at time t. Then sell them at time t+2 b) Invest $100k in stock A and $100k in stock B at time t. Then rebalance the portfolio with (0.5, 0.5) weight at time t+1. That is, buy and sell a fraction of stock A and B in your portfolio so that the dollar values of stock A and B in your portfolio are the same. Then sell them at time t+2. What would be the net return of your (two-period) investment in each case a) and b)? Hint: See LN3 Page 23. I didnt change a thing. Try to understand what happed at t+ in that page. Rebalancing means buy some assets and sell the other assets holding the total value of portfolio the same. If you compute how the dollar amount invested in each stock at each time(t 1, t+2), then you will find the total dollar value of your portfolio in the beginning (t) and in the end (t+2) at each case of a) and b). Then compute the net return. 1) Case a) → 0%, Case b) → 0% 2) Case a)-) 25%, Case b) 090 3) Case a)-) 0%, Case b)-) 25% 4) Case a)-) 56.25%, Case b)-) 0% 5) Case a)-) 0%, Case b)-) 56.25%

Example) In the above example, assume the net cash-flow C is unknown. If the discount rate changes from 0.2% to 0.1% per month, compounded monthly, what is the net return of the capital gain (or loss) of investment in the condo? Answer) 100%. PC/0.002, P2-C/0.001, r-PyP,-1 = (C/0.001) / (C/0 .002)-1-0.002/0.001-1-2-1-1-100%

15. In the above example, assume the net cash-flow C is unknown. If the discount rate changes from 0.2% to 0.5% per month, compounded monthly, what is the net return of the capital gain (or loss) of investment in the condo? (Hint l: net return (%)-100 (nua price answer will be the same regardless of the value of C. So, you can pick whatever number for C and still find the same answer. Find the old and new prices and then compute the net return.) old price 1), Hint 2: The D-10% 2)-30% 3)-50% 4)-60% 5)-80% 16. In the example above Q15, if the net cash-flow C is doubled and the discount rate r is tripled, what would be the net retum of the capital gain (or loss) of investment in the condo? (Hint: The answer will be the same regardless of the value of the old C and the original r. So, you can pick whatever numbers for C and r, and you will still find the same answer. Find the old and nevw prices and then compute the net return.) 1 )-33.33% 2)-30. 1 5% 3)-25.54% 4)-21.68% 5)-18.96%
Answer #1

Question 15

P1 = C/0.002

P2 = C/0.005

r = P2/P1 -1 =(C/0.005)/ (C/0.002 )-1 = 0.002/0.005 -1 = -0.6 = -60%

Answer : -60% (Option 4)

Question 16

P1 = C/0.002

P2 = 2C/0.006

r = P2/P1 -1 =(2C/0.006)/ (C/0.002 )-1 = 0.004/0.006 -1 = -0.3333 = -33.33% (Option 1)

Answer : -33.33% (Option 1)